# Traffic Scene Generation from Natural Language Description for Autonomous Vehicles with Large Language Model

### v1

@@ Abstract
Text-to-scene generation, transforming textual descriptions into detailed scenes, typically relies on generating key scenarios along predetermined paths, constraining environmental diversity and limiting customization flexibility. To address these limitations, we propose a novel text-to-traffic scene framework that leverages a large language model to generate diverse traffic scenarios within the Carla simulator based on natural language descriptions. Users can define specific parameters such as weather conditions, vehicle types, and road signals, while our pipeline can autonomously select the starting point and scenario details, generating scenes from scratch without relying on predetermined locations or trajectories. Furthermore, our framework supports both critical and routine traffic scenarios, enhancing its applicability. Experimental results indicate that our approach promotes diverse agent planning and road selection, enhancing the training of autonomous agents in traffic environments. Notably, our methodology has achieved a 16% reduction in average collision rates. Our work is made publicly available at https://basiclab.github.io/TTSG/.
@@ I Introduction
In autonomous systems, the ability of agents to effectively interact with their surroundings, recognize hazardous objects in the environment and perform safe maneuvers enhances the accuracy of traffic trajectory prediction and planning. To achieve these tasks, datasets such as nuScenes [1, 2] and Waymo [3, 4] are made to provide substantial real-world data, aiding researchers to train a model that can handle driving conditions under several environments and implicitly recognizing different driving rules. Nevertheless, the existing datasets often lack variability in critical conditions or scene types, and it can be hard for agents to interact with the environment since conducting diverse real-world tests with autonomous agents is costly and poses safety risks. Consequently, there is increasing reliance on simulated environments [5, 6] to gather data and evaluate agents. These simulations, however, typically require predefined routes and spawning positions for agents, which limits testing flexibility.
@@
We aim to generate traffic scenes through natural language inputs, enabling dynamic and flexible scenario creation. For instance, given an input prompt: “Schedule a daily traffic at the intersection”, we can obtain a traffic scenario that matches the description. Recent advancements have highlighted the powerful capabilities of Large Language Models (LLMs) in reasoning [7, 8], planning [9, 10], and even code generation [11]. One existing approach designed specifically for generating critical scenarios, ChatScene [12], leverages an LLM to generate critical conditions within the Carla environment [5]. It does so by first selecting a few predefined positions and agents to perform the scenario and adopting an LLM agent to perform in-context learning with a few code examples to generate code for controlling agents. However, while ChatScene uses predefined positions to reduce the complexity of generating scenarios, it sacrifices diversity. Furthermore, the reliance on few-shot examples and fixed formats restrict customization options, such as agent types, road signals, and objects, limiting the creation of tailored traffic scenarios.
Figure 1: Comparison between ChatScene (top) and ours (bottom).
@@
To increase diversity and allow more customized options, we introduce a novel pipeline that can directly provide the agent scheduling, including types, actions, positions, and road conditions, and can automatically select proper positions for scene generation without requiring human-specified trajectories or spawning positions. The comparison of ChatScene and our pipeline is shown in Fig. 1. Specifically, our system first uses the LLM agent to perform the prompt analysis to decompose different road conditions and required agents. The potential roads that match the condition are retrieved from the database. Then, our system prompts the LLM agent to provide the agent planning that should be created within the scene based on the analysis results. Since agents can take different actions, we should choose the road that allows them to perform specific actions. For instance, there should be a left road if an agent is going to turn left. Therefore, we rank the potential roads retrieved before with the agent’s plans to find the highest one for scene generation. Finally, we design our rendering interface to translate the planning results from LLM into the traffic scenario. Since our approach can automatically select the road for spawning and place the agent based on their plans, no predefined selection of points and trajectories is required. We also build an agent database that includes different types of agents and actions that enable users to choose from.
@@
We demonstrate the efficacy of our framework by training an agent on our generated critical scenarios and evaluating the generation quality under different component designs. Our contributions are summarized as follows:
@@
  • We propose a novel system for generating traffic scenes from natural language descriptions through road retrieval and agent planning pipeline with an LLM.
@@
  • Our system supports the diverse generation of traffic scenes on road selection and agent planning and allows customized options for different scenario usages.
  • Our approach can be used to train an agent under critical scenarios and obtain a 16% reduction in collision rate.
@@ II-A Traffic Scene Generation
A recent line of work focuses on critical scenario generation as it poses significant risks during driving [19, 20, 21]. These techniques often adopt adversarial training under scenarios initially generated in critical states to ensure that driving agents can manage critical cases and enhance safety. To generate the critical scenario automatically with text, ChatScene [12] facilitates text-to-critical-scenario conversion using an LLM agent that employs the Scenic programming language [22] to manage autonomous agents. Unlike previous work, we implement a road-and-agent retrieval system that facilitates text-to-scene conversions applicable to diverse scenarios, including multi-agent driving and critical scenario production. This approach allows for greater flexibility compared to ChatScene, which necessitates predefined spawning points and offers limited customized options by enabling users to generate scenes directly from text without the need to select maps or locations.
@@ II-B Autonomous Driving with LLM
Large Language Models (LLMs), such as GPT series [23] and Llama [24], have demonstrated substantial capabilities in natural language processing and have been increasingly adapted for diverse applications [25, 26, 27]. In the domain of autonomous driving, one key objective is to enable decision-making based on robust reasoning. Due to their potent planning and reasoning capabilities, LLMs have been integrated into vision language model (VLM) systems for trajectory scheduling and identifying safety-critical objects.
For instance, ADAPT [7] utilizes LLMs for video captioning and reasoning of the driving decisions. Regarding action planning, DriveLM [9] and DriveLLM [28] employ LLMs to dynamically plan vehicle actions based on image inputs. To incorporate different modalities such as point clouds, DriveMLM [10] apply Q-former [29] to translate all inputs into tokens that can be read by an LLM. To deploy the agent into a continuous driving environment, DiLu [30] incorporates a memory module within its framework, enabling the continuous learning and application of new knowledge derived from environmental interactions. Additionally, RAG-Driver [31] implements a RAG system to analogize current driving conditions with historical scenarios, thereby providing precise planning inputs with past experience.
@@ III Traffic Scene Generation
Figure 2: Illustration of the TTSG pipeline, detailing five principal stages: analysis, road candidate retrieval, agent planning, road ranking, and generation. The analysis stage decompose the user’s input to refine the processing in subsequent phases. This is followed by an LLM agent determining retrieval conditions to identify appropriate road candidates and facilitating precise planning for agents, which includes types, positions, and actions. In cases where multiple road options are retrieved, a ranking strategy prioritizes roads that best match the desired conditions and offer sufficient lanes and space for scenario depiction. The traffic scene is ultimately generated through our custom rendering interface, utilizing the selected road and agent details.
In this section, we detail the development of our Text-to-Traffic Scene Generation (TTSG) framework, which leverages an LLM for planning and retrieval to generate traffic scenes from natural language descriptions. Our methodology starts with constructing a database that categorizes the attributes of roads and agents. Given a scene description as shown in Fig. 2-(a), an LLM agent first analyzes the input prompts, determining the road conditions for retrieval and devising a plan for agent scheduling. Following this, the potential road candidates undergo a ranking process, which evaluates the properties and actions of agents to identify the most appropriate selection for scene generation. This ranking strategy is crucial for ensuring the optimal match of road properties to the intended scenario. The characteristics of the selected road are then integrated with the agents’ planning outcomes and visualized using our custom rendering interface. To validate our pipeline’s effectiveness and conduct thorough evaluations, the Carla simulator is utilized as our primary demonstration platform. For effective record--keeping and system integration, the outputs from the LLM agent are all stored in JSON format.
@@ III-A Road and Agent Database Construction
Constructing a road database is crucial for augmenting the capabilities of the large language model (LLM) agent. This database is designed to provide spawn information, enabling the LLM to deliver precise planning results even for scenarios it has not previously encountered. Unlike previous work [12], which focuses on gathering code snippets, our database categorizes roads, including traffic signals and road objects found within each road segment. This strategy allows for dynamic scenario generation without the need for predefined geometries and spawning positions, enhancing the flexibility of our system across various traffic conditions.
@@ III-B1 Analysis
The primary objective in the analysis mode is to equip the LLM agent with a complete understanding of the user’s requirements, facilitating precise planning decisions. One promising approach is to leverage the chain-of-thought method [32, 33], which involves generating step-by-step explanations. However, this approach requires more output tokens from the LLM, which can be computationally expensive. To streamline the process, we instead aim to provide the LLM agent with enriched contextual knowledge by first decomposing each component, including the required signals, objects, and agents from the user’s input as illustrated in Fig. 2-(a). This decomposed output aids the LLM in processing and interpreting the user’s natural language description more effectively, setting the stage for a precise understanding of the input for scenario generation.
@@ III-B2 Road Candidate Retrieval
This mode focuses on querying the road database to find the road that aligns with the user’s specifications and ensures accurate scene generation. The retrieval process is crucial for selecting optimal road locations and obtaining essential information for scenario planning. Leveraging the output from the analysis phase and the user’s prompt, our system efficiently narrows the search to identify roads within the road database that meet specific criteria, including the number of lanes, having or not having any specific traffic signals, and the roadside objects (e.g., speed signs or stop signs).
@@ III-C2 Critical Scenario
A significant aspect of our scene generation capability focuses on critical scenarios, which are essential for testing the robustness of autonomous driving systems. Our system can be prompted to create scenarios with potentially dangerous settings. For example, in Fig. 4(a), we can generate scenes where visibility is obstructed given input presented in the sub-caption, causing pedestrians and vehicles to be unseen by each other. Additionally, with the road conditions, we can create an unprotected left-turn scene with no traffic light and stop sign as presented in Fig. 4(b).
@@ IV-A Pipeline Setup
In the experimental setup, we utilize the Carla simulator, version 0.9.15, and employ the built-in maps from Town01 to Town07 to construct our road database. Roads lacking sufficient space for agent deployment are excluded. Our agent database accommodates nine distinct types of agents: ambulance, police car, firetruck, bus, truck, motorcycle, cyclist, car, and pedestrian. The simulation engine supports six actions: turning, lane changing, straight, stop, crossing road, and blocking, with denoting optional left or right directions. Additionally, as for position relative to the ego agent, our system supports twelve directional positions, including eight basic compass directions and specific placements like left road, right road, opposite straight road, and near the ego destination. We provide the illustration for the action and the relative position in Fig. 5. As for the LLM agent, we adopt the GPT-4o [23] to achieve analysis, road candidate retrieval, and agent planning.
@@ IV-B Generation for Critical Scenarios
To assess the efficacy of our framework to train an autonomous agent, we examine our approach to critical scenarios and utilize SafeBench [34], a benchmark designed for the safety evaluation of autonomous vehicle systems. Our pipeline is employed to create scenes across three distinct challenging safety scenarios selected from ChatScene [12], aiming to conduct adversarial training to bolster the ego vehicle’s safe-driving skills. We prompt the LLM agent with the scenario descriptions and use the planning results from the LLM and our rendering interface to train the ego agents under a controlled policy using a soft-actor-critic model [35]. During training, the ego agent is controlled by a learnable policy model, and other agents that create the scenario are controlled by our engine.
For consistency and comparability with the studies reported in [12], the vehicles are evaluated in the same environment of [12] to measure their Collision Rate (CR) and Overall Score (OS), which includes various metrics of driving performance. The outcomes are presented in Table I, and each column is colored as best and second best.
In this table, we compare our results with Learning-to-Collide (LS) [36], AdvSim (AS) [37], Adversarial Trajectory Optimization (AT) [38], and ChatScene (CS) [12]. Our findings indicate that our text-to-scene pipeline effectively trains agents to handle critical conditions, thereby enhancing safety and reducing collision risks. Notably, our agents are proficient at managing obstacles and yielding to oncoming cars, thus avoiding dangerous collisions in the first and third scenario types. Despite achieving a higher collision rate and a lower driving score in the second scenario, where agents occasionally fail to decelerate when another vehicle attempts a risky maneuver, our system demonstrates competence in most cases to deal with the dangerous lane-changing scenarios, surpassing the majority of comparative approaches. Overall, our approach successfully prepares agents for critical scenarios, enabling them to learn from these experiences and elevate their safety capabilities.
@@
TABLE I: Results evaluated on Critical Scenario.
Algo. Metric Scenario Avg.
Straight Obstacle Lane Changing Unprotected Left-turn
[…]
AS 0.230 0.530 0.050 0.270
AT 0.140 0.300 0.000 0.150
CS 0.030 0.110 0.100 0.080
Ours 0.000 0.200 0.000 0.067
[…]
AS 0.784 0.666 0.937 0.796
AT 0.849 0.803 0.948 0.867
CS 0.905 0.906 0.903 0.905
ours 0.955 0.861 0.951 0.922

### v2

@@ Abstract
Text-to-scene generation typically limits environmental diversity by generating key scenarios along predetermined paths. To address these constraints, we propose a novel text-to-traffic scene framework that leverages a large language model (LLM) to autonomously generate diverse traffic scenarios for the CARLA simulator based on natural language descriptions. Our pipeline comprises several key stages: (1) Prompt Analysis, where natural language inputs are decomposed; (2) Road Retrieval, selecting optimal roads from a database; (3) Agent Planning, detailing agent types and behaviors; (4) Road Ranking, scoring roads to match scenario requirements; and (5) Scene Generation, rendering the planned scenarios in the simulator. This framework supports both routine and critical traffic scenarios, enhancing its applicability. We demonstrate that our approach not only diversifies agent planning and road selection but also significantly reduces the average collision rate from 8% to 3.5% in SafeBench. Additionally, our framework improves narration and reasoning for driving captioning tasks. Our contributions and resources are publicly available at https://basiclab.github.io/TTSG/.
@@ I Introduction
In autonomous systems, motion planning models often require a lot of data training to earn robust driving capabilities. An interactive training environment that can generate scenarios based on instructions can greatly accelerate model training and reduce reliance on large amounts of data. Previous work on datasets, such as nuScenes [1, 2] and Waymo [3, 4] are made to provide substantial real-world data, aiding researchers to train a model that can handle driving conditions under several environments and implicitly recognizing different driving rules. Nevertheless, the existing datasets often lack variability in critical conditions or scene types, and it can be hard for agents to interact with the environment since conducting diverse real-world tests with autonomous agents is costly and poses safety risks. Consequently, there is increasing reliance on simulated environments [5, 6] to gather data and evaluate agents. These simulations, however, typically require predefined routes and spawning positions for agents, which limits testing flexibility.
@@
We aim to generate traffic scenes through natural language inputs, enabling dynamic and flexible scenario creation. For instance, given an input prompt: “Schedule a daily traffic at the intersection”, we can obtain a traffic scenario that matches the description. Recent advancements have highlighted the powerful capabilities of Large Language Models (LLMs) in reasoning [7, 8], planning [9, 10], and even code generation [11]. A previous approach, ChatScene [12], has proposed to generate specific critical scenarios in CARLA [5]. However, as [12] uses predefined positions to reduce the complexity of generating scenarios, it sacrifices diversity. Furthermore, the reliance on few-shot examples and fixed formats restrict customization options, such as agent types, road signals, and objects, limiting the creation of tailored traffic scenarios.
Figure 1: Simplified pipeline of our approach. Our approach takes the scene description from the users and provides the correspondence traffic scene.
@@
In this work, we introduce a novel pipeline that can directly provide the agent scheduling, including types, actions, positions, and road conditions, and can automatically select proper positions for scene generation without requiring human-specified spawning positions. The simplified pipeline of our framework is shown in Fig. 1. Specifically, our framework first uses the LLM agent to perform the prompt analysis to decompose different road conditions and required agents. The potential roads that match the condition are retrieved from the database. Then, our framework prompts the LLM agent to provide the agent planning that should be created within the scene based on the analysis results. Since agents can take different actions, the pipeline should choose the road that allows the agents to perform specific actions. For instance, there should be a left road if an agent is going to turn left. Therefore, we insert a ranking mechanism in the pipeline to score potential roads by counting the achieved requirements of the agent’s plans to find the highest one for scene generation. Finally, we design our rendering interface to translate the planning results from LLM into the traffic scenario. Since our approach can automatically select the road for spawning and place the agent based on their plans, no predefined selection of points and maps is required. We also build an agent database that includes different types of agents and actions that enable users to choose from.
@@
We demonstrate the efficacy of our framework by training an agent on our generated critical scenarios for both navigating and captioning tasks. To study the importance of each component, we evaluate the generation quality under different designs to provide a deeper understanding of our pipeline. Our contributions are summarized as follows:
@@
  • We propose a novel framework for generating traffic scenes from natural language descriptions through road retrieval and agent planning pipeline with an LLM.
@@
  • Our framework supports the diverse generation of traffic scenes on road selection and agent planning and allows customized options for different scenario usages.
  • Our approach can be used to train an agent under critical scenarios to reduce collision rate and to improve driving score and can also be incorporated to generate data for training a captioning model to narrate and reason the actions for critical scenarios.
@@ II-A Traffic Scene Generation
A recent line of work focuses on critical scenario generation as it poses significant risks during driving [19, 20, 21]. These techniques often adopt adversarial training under scenarios initially generated in critical states to ensure that driving agents can manage critical cases and enhance safety. To generate the critical scenario automatically with text, ChatScene [12] facilitates text-to-critical-scenario conversion using an LLM agent that employs the Scenic programming language [22] to manage autonomous agents. Unlike previous work, we implement a road-and-agent retrieval method that facilitates text-to-scene conversions applicable to diverse scenarios, including multi-agent driving and critical scenario production. This approach allows for greater flexibility without necessitating predefined spawning points and offers various customized options by enabling users to generate scenes directly from text.
@@ II-B Autonomous Driving with LLM
Large Language Models (LLMs), such as GPT series [23] and Llama [24], have demonstrated substantial capabilities in natural language processing and have been increasingly adapted for diverse applications [25, 26, 27, 28]. In the domain of autonomous driving, one key objective is to enable decision-making based on robust reasoning. Due to their potent planning and reasoning capabilities, LLMs have been integrated into vision language models (VLMs) for trajectory scheduling and identifying safety-critical objects.
For instance, ADAPT [7], and DriveGPT4 [29] utilize LLMs for video captioning and reasoning of the driving decisions. Regarding action planning, DriveLM [9] and DriveLLM [30] employ LLMs to dynamically plan vehicle actions based on image inputs. To incorporate different modalities such as point clouds, DriveMLM [10] apply Q-former [31] to translate all inputs into tokens that can be read by an LLM. To deploy the agent into a continuous driving environment, DiLu [32] incorporates a memory module within its framework, enabling the continuous learning and application of new knowledge derived from environmental interactions. Additionally, RAG-Driver [33] implements a RAG system to analogize current driving conditions with historical scenarios, thereby providing precise planning inputs with past experience.
@@ III Traffic Scene Generation
Figure 2: Illustration of the TTSG pipeline, detailing five principal stages: analysis, road candidate retrieval, agent planning, road ranking, and generation. The analysis stage decomposes the user’s input to refine the processing in subsequent phases. This is followed by an LLM agent determining retrieval conditions to identify appropriate road candidates and facilitating precise planning for agents, which includes types, positions, and actions. In cases where multiple road options are retrieved, a ranking strategy prioritizes roads that best match the desired conditions and offer sufficient lanes and space for scenario depiction. The traffic scene is ultimately generated through our custom rendering interface, utilizing the selected road and agent details.
In this section, we detail the development of our Text-to-Traffic Scene Generation (TTSG) framework, which leverages an LLM for planning and retrieval to generate traffic scenes from natural language descriptions. Our methodology starts with constructing a database that categorizes the attributes of roads and agents (Sec. III-A). Given a scene description as shown in Fig. 2(a), an LLM agent first analyzes the input prompts, determining the road conditions for retrieval and devising a plan for agent scheduling (Sec. III-B1, III-B2 and III-B3). Following this, the potential road candidates undergo a ranking process, which evaluates the properties and actions of agents to identify the most appropriate selection for scene generation (Sec. III-B4). This ranking strategy is crucial for ensuring the optimal match of road properties to the intended scenario. The characteristics of the selected road are then integrated with the agents’ planning outcomes and visualized using our custom rendering interface. To validate our pipeline’s effectiveness and conduct thorough evaluations, the CARLA simulator is utilized as our primary demonstration platform. For effective record--keeping and system integration, the outputs from the LLM agent are all stored in JSON format.
@@ III-A Road and Agent Database Construction
Constructing a road database is crucial for augmenting the capabilities of the large language model (LLM) agent. This database is designed to provide spawn information, enabling the LLM to deliver precise planning results even for scenarios it has not previously encountered. Unlike previous work [12], which focuses on gathering code snippets, our database categorizes roads, including traffic signals and road objects found within each road segment. This strategy allows for dynamic scenario generation without the need for predefined geometries and spawning positions, enhancing the flexibility of our framework across various traffic conditions.
@@ III-B1 Analysis
The primary objective in the analysis mode is to equip the LLM agent with a complete understanding of the user’s requirements, facilitating precise planning decisions. One promising approach is to leverage the chain-of-thought method [34, 35], which involves generating step-by-step explanations. However, this approach requires more output tokens from the LLM, which can be computationally expensive. To streamline the process, we instead aim to provide the LLM agent with enriched contextual knowledge by first decomposing each component, including the required signals, objects, and agents from the user’s input as illustrated in Fig. 2(a). This decomposed output aids the LLM in processing and interpreting the user’s natural language description more effectively, setting the stage for a precise understanding of the input for scenario generation.
@@ III-B2 Road Candidate Retrieval
This mode focuses on querying the road database to find the road that aligns with the user’s specifications and ensures accurate scene generation. The retrieval process is crucial for selecting optimal road locations and obtaining essential information for scenario planning. Leveraging the output from the analysis phase and the user’s prompt, our framework efficiently narrows the search to identify roads within the road database that meet specific criteria, including the number of lanes, having or not having any specific traffic signals, and the roadside objects (e.g., speed signs or stop signs).
@@ III-B5 Scene Generation
TABLE I: Results evaluated on Critical Scenario with SafeBench.
[…]
Straight Obstacle Lane Changing Unprotected Left-turn Average Straight Obstacle Lane Changing Unprotected Left-turn Average
LC [36] IROS’20 0.120 0.510 0.000 0.210 0.827 0.684 0.954 0.822
AS [37] CVPR’21 0.230 0.430 0.050 0.270 0.784 0.666 0.937 0.796
AT [38] CVPR’22 0.140 0.300 0.000 0.150 0.849 0.803 0.948 0.867
CS [12] CVPR’24 0.030 0.110 0.100 0.080 0.905 0.906 0.903 0.905
ours 0.021 0.085 0.000 0.035 0.895 0.894 0.953 0.914
@@ III-C2 Critical Scenario
A significant aspect of our scene generation capability focuses on critical scenarios, which are essential for testing the robustness of autonomous driving systems. Our framework can be prompted to create scenarios with potentially dangerous settings. For example, in Fig. 4(a), we can generate scenes where visibility is obstructed given input presented in the sub-caption, causing pedestrians and vehicles to be unseen by each other. Additionally, with the road conditions, we can create an unprotected left-turn scene with no traffic light and stop sign as presented in Fig. 4(b).
@@ IV-A Pipeline Setup
In the experimental setup, we utilize the CARLA simulator and the built-in maps from Town01 to Town07 to construct our road database. Roads lacking sufficient space for agent deployment are excluded. Our agent database accommodates nine distinct types of agents: ambulance, police car, firetruck, bus, truck, motorcycle, cyclist, car, and pedestrian. The simulation engine supports six actions: turning, lane changing, straight, stop, crossing road, and blocking, with denoting optional left or right directions. Additionally, as for position relative to the ego agent, our framework supports twelve directional positions, including eight basic compass directions and specific placements like left road, right road, opposite straight road, and near the ego destination. We illustrate the action and the relative position in Fig. 5. As for the LLM agent, we adopt the GPT-4o [23] to achieve analysis, road candidate retrieval, and agent planning.
@@ IV-B Generation for Critical Scenarios
To assess the efficacy of our framework to train an autonomous agent, we examine our approach to critical scenarios with SafeBench [39], a benchmark for the safety evaluation of autonomous driving systems. Our pipeline creates scenes across three distinct challenging safety scenarios selected from ChatScene [12], aiming to conduct adversarial training to bolster the ego vehicle’s safe-driving skills. We prompt the LLM agent with the scenario descriptions and use the planning results from the LLM and our rendering interface to train the ego agents under a controlled policy using a soft-actor-critic model [40]. During training, the ego agent is controlled by a learnable policy model, and our engine controls other agents that create the scenario.
[…]
In this table, we compare our results with Learning-to-Collide (LC) [36], AdvSim (AS) [37], Adversarial Trajectory Optimization (AT) [38], and ChatScene (CS) [12]. Our findings indicate that our text-to-scene pipeline effectively trains agents to handle critical conditions, thereby enhancing safety and reducing collision risks. Notably, since our pipeline can provide diverse scenes during training, our agents are proficient at managing obstacles, lane changing, and yielding to oncoming cars. This can be observed by reducing the average collision rate from 0.08 to 0.035. However, reducing collisions may come at the cost of comfort (e.g., sudden swerving or hard braking), leading to slightly lower overall scores in certain scenarios. Overall, our approach effectively enables agents to learn from critical scenario experiences and enhances their safety capabilities.